18) Evaluate ∮(z)=1 Z² Sin1/z dZ?


1
Expert's answer
2021-09-25T11:36:32-0400

On the unit circle,"z = sin(iθ)\space and \space dz = ie^{iθ} dθ" . We then have


"\oint|z|=1 z + 1 z2 dz = Z\intop^ {2π} _0 (e−iθ+e−2iθ)ie^{iθ} dθ"

"= i Z\intop^ {2π} _0 (1+e−iθ) dθ = 2πi"


The integrand "(z + 1)\/z^2" has a double pole at z = 0. The Laurent expansion in a deleted neighborhood of z = 0 is simply "1( z + 1) z^2" , where the coefficient of 1/z is seen to be 1. We have 


"Res (\frac{z+1}{z^2},0)= 1"